# Efficient Implementation of Gaussian Process Regression Accelerated Saddle Point Searches with Application to Molecular Reactions

### v1

@@ Abstract
The task of locating first order saddle points on high-dimensional surfaces describing the variation of energy as a function of atomic coordinates is an essential step for identifying the mechanism and estimating the rate of thermally activated events within the harmonic approximation of transition state theory. When combined directly with electronic structure calculations, the number of energy and atomic force evaluations needed for convergence is a primary issue. Here, we describe an efficient implementation of Gaussian process regression (GPR) acceleration of the minimum mode following method where a dimer is used to estimate the lowest eigenmode of the Hessian. A surrogate energy surface is constructed and updated after each electronic structure calculation. The method is applied to a test set of 500 molecular reactions previously generated by Hermez and coworkers [J. Chem. Theory Comput. 18, 6974 (2022)]. An order of magnitude reduction in the number of electronic structure calculations needed to reach the saddle point configurations is obtained by using the GPR compared to the dimer method. Despite the wide range in stiffness of the molecular degrees of freedom, the calculations are carried out using Cartesian coordinates and are found to require similar number of electronic structure calculations as an elaborate internal coordinate method implemented in the Sella software package. The present implementation of the GPR surrogate model in C++ is efficient enough for the wall time of the saddle point searches to be reduced in 3 out of 4 cases even though the calculations are carried out at a low Hartree-Fock level.
@@ 1 Introduction
An important task in the modelling of thermally induced transitions, such as chemical reactions, diffusion events and conformational changes of molecules, is finding the mechanism and estimating the corresponding rate. For transitions that are slow on the time scale of atomic vibrations, the rate can be estimated using transition state theory and when the atoms are in a confined environment vibrating about a well defined average position the harmonic approximation can suffice. Then, the key challenge is finding first-order saddle points on the energy surface describing the variation of the system with respect to the atomic coordinates. At a first-order saddle point, the atomic forces vanish and the Hessian matrix has a single negative eigenvalue. Several methods to systematically identify saddle points have been developed over the past decades, with varying implementations in software packages (see, for example [36, 14]). When combined directly with electronic structure calculations, the key issue is to reduce as much as possible the number of times the energy and atomic forces need to be calculated as each calculation may require hours of computations.
@@
There are two types of saddle point searches. Either both the initial state minimum and the final state minimum are specified beforehand, i.e. a a two endpoint boundary condition, or only the initial state is specified beforehand, i.e. a single endpoint boundary condition. The former is often approached by discretising the path and converging on a minimum energy path connecting the two endpoints, using for example, the nudged elastic band (NEB) method [22, 2]. The latter is more challenging and can provide unexpected results for the transition mechanism as well as the resulting final state. Such calculations can, therefore be used as the basis for simulations of the long timescale evolution of a system with, e.g. the adaptive kinetic Monte Carlo method [16]. Furthermore, single endpoint methods can provide an efficient way to complete two endpoint calculations when a rough convergence to the saddle point has been obtained [2].
@@
Machine learning can be used to accelerate searches for saddle points in many ways. A potential energy surface can be generated to mimic the atomic interactions in a system using various machine learning approaches with input from electronic structure calculations and then the evaluation of energy and atomic forces in saddle point searches are relatively fast. The accuracy of such energy surfaces is, however, often not good enough in regions near saddle points because training sets tend to lack data from these regions. It is then important to keep retraining the enery surface as information on saddle point regions is acquired. The training of a potential energy surface in this manner requires a large set of electronic structure calculations. When many systems are being screend for a particular property, using for example, workflow engines (e.g. AiiDA [20], Snakemake [30], etc.), the training of a transferable potential enery surface for each candidate is not viable because it requires a large investment in training set selection. It is then better to use machine learning to accelerate each saddle point search using just the minimum number of electronic structure calculations required to reach convergence. This is the approach demonstrated here, namely the acceleration of a single saddle point search so as to reduce the number of electronic structure calculations as much as possible.
@@
Cartesian coordinates have most often been used in saddle point search algorithms because of the ease of implementation [35, 19, 24, 32]. However, the various degrees of freedom of a molecule, for example, can have a wide range in stiffness, for example, vibrations of strong covalent bonds vs. rotation of a methyl group. It can, therefore, be advantageous to use “internal coordinates” instead of Cartesian coordinates [41]. The automated construction of these internal coordinates can be challenging when considering a wide range of atomic structures. This is particularly a challenge when near linear configurations of three or more atoms are present. In calculations of heterogeneous catalysis involving the reaction of molecules and surfaces of crystals, such linear arrangements of atoms occurs frequently. Recently, an elaborate saddle point search method based on nonredundant internal coordinates has been presented and implemented in the software called Sella [18]. The algorithm involves a geodesic update of constrained internal reaction coordinates [17]. When near-linearity in the arrangement of atoms is encountered, ghost atoms are introduced in an automated manner to help define meaningful internal coordinates. The efficacy of the method has been demonstrated in calculations of 500 systems generated using a database approach for predicting saddle point configurations of the atoms. We note that this data set contains only systems where the saddle point searches using Sella converge [18].
@@
Here, as noted earlier, we present a different approach, namely the use of Cartesian coordinates in combination with a surrogate energy surface using with a GPR. A local approximation to the energy surface is generated using the electronic structure calculations carried out during each saddle point search. The eigenvector corresponding to the lowest eigenvalue of the Hessian, the so-called minimum mode, is used to define the search direction and it is found using the dimer method [15]. We refer to this method as the GPR-dimer. An earlier version of this algorithm has been presented [26], but the work presented here uses an improved implementation in C++ within the EON software package [5]. We directly compare the performance of our method to the internal coordinate approach of Sella and the dimer method for the 500 system dataset. Our results show that by using the GPR-dimer, the number of electronic structure calculations needed to converge to saddle points can be reduced by an order of magnitude compared to the dimer and the overall computational time can be reduced in most cases by using the GPR despite the additional overhead even though the electronic structure calculations are carried out at a low Hartree-Fock (HF) level. A performance similar to that of Sella is obtained even though Cartesian coordinates are used for this dataset of molecular reactions. In some cases, the GPR dimer method convreges on saddle points that are closer to the initial guess, both in terms of atomic coordinates and total energy.
@@
The article is organized as follows: In Section 2, the saddle point search method and GPR is reviewed for completeness. The performance of saddle point searches for the 500 systems is presented in Section 3, and in 4, a discussion of results is presented. We conclude in Section 5.
@@ 2.1 Minimum mode following with a dimer
The minimum mode following (MMF) method [15, 38] is a technique for finding first order saddle points on a surface by startin from a point that is displaced from the initial state energy minimum. The search is subsequently guided by the eigenvector corresponding to the lowest eigenvalue of the Hessian, i.e. the minimum mode [15]. If at least one of the eigenvalues of the Hessian is negative, the component of the atomic force in the direction of the minimum mode is inverted to form the search direction. Otherwise the search direction is uphill along the minimum mode. The minimum mode can be found in several ways without explictly constructing the Hessian in memory. One option is to use a “dimer” which consists of two replicas of the system (i.e. confgiruations of all atoms) separated by a small, fixed distance, $\Delta R$. Letting the two images be denoted as $\mathbf{R}_{1}$ and $\mathbf{R}_{2}$, they are positioned on eithe side of the midpoint $\mathbf{R}$, such that $\mathbf{R}_{1}=\mathbf{R}+\Delta R\hat{\mathbf{N}}$ and $\mathbf{R}_{2}=\mathbf{R}-\Delta R\hat{\mathbf{N}}$, where $\hat{\mathbf{N}}$ is a unit vector defining the dimer’s orientation. Starting from an initial guess for the midpoint $\mathbf{R}$, for example using a preconceived approximate saddle point configuration, and a randomly oriented unit vector $\hat{\mathbf{N}}$, the energy of each image, ($E_{1}$ and $E_{2}$), with the correspondin atomic forces ($\mathbf{F}_{1}$, $\mathbf{F}_{2}$) are evaluated at each image. The dimer is rotated so as to minimize the total dimer energy, $E$, which is the sum of the image energies: $E=E_{1}+E_{2}$ while keeping the separation between the images fixed. This aligns the orientation of the dimer, $\hat{\mathbf{N}}$, with the minimum mode. An iterative process is used to converge to the direction of the minimum mode to a given tolerance in the force [35]. During the rotation phase, the atomic forces are projected to make them perpendicular to the current dimer orientation, $\mathbf{F}^{\perp}$. Each rotational step involves finding a plane defined by the rotational force and the previous rotation plane, and the rotational force is then zeroed within this plane. Here, a conjugate gradient (CG) method is used where at each step a linear combination of the force vector and the previous search direction is calculated. The coefficient of the previous direction, or the CG parameter, is also a sensitive indicator of convergence. The Polak-Ribiere method [39]:
@@
Once the dimer has been oriented sufficiently close to the minimum mode, a translational step moves the midpoint of the dimer forward. If at least one of the eigenvalues of the Hessian is negative, a modified force, $\mathbf{F}^{\dagger}$, is used, where the component of the force along the dimer direction is inverted: $\mathbf{F}_{R}-2(\mathbf{F}_{R}\cdot\hat{\mathbf{N}})\hat{\mathbf{N}}$. This inversion ensures that the dimer moves uphill on the energy surface along the minimum mode and downhill along all other directions. The dimer is iteratively rotated and translated in this way until the atomic forces become zero to a given tolerance, indicating that a saddle point has been located. The limited memory Broyden-Fletcher-Goldfarb-Shanno algorithm (L-BFGS) on the midpoint of the dimer at each translation step, as it has previously been found to be efficient [24, 42].
@@ 2.2 Gaussian Process Regression
A Gaussian Process Regression (GPR) is a non-parametric method for learning from data. It can be understood as a generalization of multivariate Gaussian distributions to function spaces, effectively providing a distribution over possible functions that fit the observed data. Here, we use a GPR to learn a relationship between Cartesian coordinates of the atoms and their corresponding energy and atomic forces. The GPR model infers this relationship by finding similarities, determined by means of a covariance function (often referred to as the kernel). The expressivity of these models stems from the fact that they may be understood as the inifinite limit of a neural network [33] and that any neural network of a finite width is essentially a GP [44].
@@
where the summations are over all atoms in the system, $(\mathbf{x},\mathbf{x}^{\prime})$ are the input vectors of atomic coordinates, $(r_{i,j}(\mathbf{x})=\sqrt{\sum_{d=1}^{3}(x_{i,d}-x_{j,d})^{2}})$ is the Euclidean distance between atoms $i$ and $j$ within configuration $\mathbf{x}$. $(\phi(i,j))$ is a function specifying the type of atom pair (i, j) and $(l_{\phi(i,j)})$ is the corresponding length scale parameter controlling how quickly the covariance changes as a function of the inverse distance between atoms. The parameter $(\sigma_{c}^{2})$ is a constant variance term while $(\sigma_{m}^{2})$ is a variance scaling factor for the distance-dependent part of the kernel.
@@
The fitting process involves the optimization of the kernel hyperparameters. The computations involve a the inversion of a matrix of dimensionality $O(M^{3}N^{3})$ and requires memory on the order of $O(M^{2}N^{2})$, where $M$ is the number of electronic structure calculations that are being used as the training set (i.e. “observations”) and $N$ is the number of atomic coordinates, three times the number of atoms. The hyperparameters are optimized using the scaled conjugate gradient method [31]. A Cholesky decomposition is used to speed up the prediction of multiple data points [40].
@@ 2.3 Computational specifications
The systems analyzed consist of 500 initial configurations of small organic molecules with between 7 and 25 atoms. The electronic structure calculations are carried out at the HF level with a 3-21G basis set using the NWChem [1] software. The spin unrestricted formalism is used for doublets and spin restricted closed shell formalism is used for the singlet configurations. The self-consistent field (SCF) threshold is set at $81e-$ 8 Hartree, and the criterion for convergence to saddle points is that the norm of the atomic force vector drops below 0.01 eV/Å. The electronic structure calculations carried out here are at the same level and use the same software as those of hermesSellaOpenSourceAutomationFriendly2022 [18].
For the GPR-dimer calcuations, which are carried out using EON [5], the geometry verification pre-processing of NWChem is turned off. For both EON and Sella calculations, the ASE [29] file I/O interface to NWChem is used, which means no symmetrization or centering takes place within NWChem, unlike the default settings for NWChem. As mentioned above, the conjugate gradient method [19] is used for the dimer rotations and the limited memory Broyden-Fletcher-Goldfarb-Shanno algorithm [34, 24] is used for the translation. Since the validity of the surrogate energy surfaces is limited to the region where data is available, early stopping criteria are used to prevent the calculation from venturing too far outside the region [26, 25].
@@
The calculations are carried out on the IRHPC elja computer cluster with snakemake as the workflow engine [30]. The cluster nodes have Intel Xeon Platinum 8358 (128 @ 2.600 GHz) CPUs which run Rocky Linux release 8.10. The Hartree Fock calculations are run with NWChem built with spack, and each run requests 16 cores exclusively for NWChem with 3 GB of memory per node. To reduce I/O overheads, the scratch directory was used. dvc was used to vendor intermediate data, and the final csv data required for the plots along with the orgmode literate programming document [11] is part of the associated mono-repo.

### v2

@@ Abstract
The task of locating first order saddle points on high-dimensional surfaces describing the variation of energy as a function of atomic coordinates is an essential step for identifying the mechanism and estimating the rate of thermally activated events within the harmonic approximation of transition state theory. When combined directly with electronic structure calculations, the number of energy and atomic force evaluations needed for convergence is a primary issue. Here, we describe an efficient implementation of Gaussian process regression (GPR) acceleration of the minimum mode following method where a dimer is used to estimate the lowest eigenmode of the Hessian. A surrogate energy surface is constructed and updated after each electronic structure calculation. The method is applied to a test set of 500 molecular reactions previously generated by Hermes and coworkers [J. Chem. Theory Comput. 18, 6974 (2022)]. An order of magnitude reduction in the number of electronic structure calculations needed to reach the saddle point configurations is obtained by using the GPR compared to the dimer method. Despite the wide range in stiffness of the molecular degrees of freedom, the calculations are carried out using Cartesian coordinates and are found to require similar number of electronic structure calculations as an elaborate internal coordinate method implemented in the Sella software package. The present implementation of the GPR surrogate model in C++ is efficient enough for the wall time of the saddle point searches to be reduced in 3 out of 4 cases even though the calculations are carried out at a low Hartree-Fock level.
@@ 1 Introduction
An important task in the modelling of thermally induced transitions, such as chemical reactions, diffusion events and conformational changes of molecules, is finding the mechanism and estimating the corresponding rate. For transitions that are slow on the time scale of atomic vibrations, the rate can be estimated using transition state theory and when the atoms are in a confined environment vibrating about a well defined average position the harmonic approximation can suffice. Then, the key challenge is finding first-order saddle points on the energy surface describing the variation of the system with respect to the atomic coordinates. At a first-order saddle point, the atomic forces vanish and the Hessian matrix has a single negative eigenvalue. Several methods to systematically identify saddle points have been developed over the past decades, with varying implementations in software packages (see, for example, Refs. 41, 18, 2, 21). When combined directly with electronic structure calculations, the key issue is to reduce as much as possible the number of times the energy and atomic forces need to be calculated as each calculation may require hours of computations.
@@
There are two types of saddle point searches. Either both the initial state and the final state minima are specified beforehand, i.e. a two endpoint boundary condition, or only the initial state is specified beforehand, i.e. a single endpoint boundary condition. The former is often approached by discretising the path and converging on a minimum energy path connecting the two endpoints, using for example, the nudged elastic band (NEB) method [27, 3]. The latter is more challenging and can provide unexpected results for the transition mechanism as well as the resulting final state. Such calculations can, therefore be used as the basis for simulations of the long timescale evolution of a system with, e.g. the adaptive kinetic Monte Carlo method [20]. Furthermore, single endpoint methods can provide an efficient way to complete two endpoint calculations when a rough convergence to the saddle point has been obtained [3].
@@
Machine learning can be used to accelerate searches for saddle points in many ways. A potential energy surface can be generated to mimic the atomic interactions in a system using various machine learning approaches with input from electronic structure calculations and then the evaluation of energy and atomic forces in saddle point searches is relatively fast. The accuracy of such energy surfaces is, however, often not good enough in regions near saddle points because training sets tend to lack data from these regions. It is then important to keep retraining the energy surface as information on saddle point regions is acquired. The training of a potential energy surface in this manner requires a large set of electronic structure calculations. When many systems are being screened for a particular property, using for example, workflow engines (e.g. AiiDA [25], Snakemake [35], etc.), the training of a transferable potential energy surface for each candidate is not viable because it requires a large investment in the training set. It is then better to use machine learning to accelerate each saddle point search using just the minimum number of electronic structure calculations required to reach convergence. This is the approach demonstrated here, namely the acceleration of a single saddle point search so as to reduce the number of electronic structure calculations as much as possible.
@@
Cartesian coordinates have most often been used in saddle point search algorithms because of the ease of implementation [40, 24, 29, 37]. However, the various degrees of freedom of a molecule can have a wide range in stiffness, for example, vibration of a strong covalent bond vs. rotation of a methyl group. It can, therefore, be advantageous to use “internal coordinates” instead of Cartesian coordinates [46]. An automated construction of internal coordinates can be challenging when considering a wide range of atomic structures. This is in particular the case when near linear configurations of three or more atoms are present. In calculations of heterogeneous catalysis involving the reaction of molecules and surfaces of crystals, such linear arrangements of atoms occur frequently. Recently, an elaborate saddle point search method based on nonredundant internal coordinates has been presented and implemented in software called Sella [23]. The algorithm involves a geodesic update of constrained internal reaction coordinates [22]. When near-linearity in the arrangement of atoms is encountered, ghost atoms are introduced in an automated manner to help define meaningful internal coordinates. The efficacy of the method has been demonstrated in calculations of 500 systems generated using a database approach for predicting saddle point configurations of the atoms. We note that this data set contains only systems where the saddle point searches using Sella converge [23].
@@
Here, as noted earlier, we present a different approach, namely the use of Cartesian coordinates in combination with a surrogate energy surface generated with GPR. A local approximation to the energy surface is generated using the electronic structure calculations carried out during each saddle point search. The eigenvector corresponding to the lowest eigenvalue of the Hessian, the so-called minimum mode, is used to define the search direction and it is found using the dimer method [19]. We refer to this method as the GPR-dimer. An earlier version of this algorithm has been presented [31], but the work presented here uses an improved implementation in C++ within the EON software package [6]. We compare the performance of our method to regular dimer calculations without a surrogate surface, as well as the internal coordinate approach of Sella for the 500 system dataset. Our results show that by using the GPR, the number of electronic structure calculations needed to converge to saddle points can be reduced by an order of magnitude and the overall computational time can be reduced in most cases despite the additional overhead even though the electronic structure calculations are carried out at a low Hartree-Fock (HF) level. A performance similar to that of Sella is found even though Cartesian coordinates are used for this dataset of molecular reactions. This is a promising result because our goal is to use GPR-dimer in calculations of molecular reactions on surfaces of crystals, where near linear configurations of atoms are common and the use of internal coordinates would require the introduction of a number of ghost atoms. In some cases, the GPR dimer method converges on saddle points that are closer to the initial guess, both in terms of atomic coordinates and total energy.
@@
The article is organized as follows: In Section 2, the saddle point search method and GPR is reviewed for completeness. The performance of saddle point searches for the 500 systems is presented in Section 3, and in Section 4, the results are discussed. Conclusions are given in Section 5.
@@ 2.1 Minimum mode following with a dimer
The minimum mode following (MMF) method [19, 43] is a technique for iteratively moving from some starting configuration of the atoms to a configuration corresponding to a first order saddle point on the energy surface. The search is guided by the eigenvector corresponding to the lowest eigenvalue of the Hessian, i.e. the minimum mode [19]. If at least one of the eigenvalues of the Hessian is negative, the component of the atomic force in the direction of the minimum mode is inverted to form the search direction. Otherwise the search direction is uphill along the minimum mode. The minimum mode can be found in several ways without explicitly constructing the Hessian.
One option is to use a “dimer” which consists of two replicas of the system (i.e. configurations of all atoms) separated by a small, fixed distance. Letting the two images be denoted as $\mathbf{R}_{1}$ and $\mathbf{R}_{2}$, they are positioned on either side of the midpoint $\mathbf{R}$, such that $\mathbf{R}_{1}=\mathbf{R}+\Delta R\hat{\mathbf{N}}$ and $\mathbf{R}_{2}=\mathbf{R}-\Delta R\hat{\mathbf{N}}$, where $\hat{\mathbf{N}}$ is a unit vector defining the dimer’s orientation and $\Delta R$ is the distance parameter. Starting from an initial guess for the midpoint $\mathbf{R}$, for example using a preconceived approximate saddle point configuration, and a randomly oriented unit vector $\hat{\mathbf{N}}$, the energy of each image, ($E_{1}$ and $E_{2}$), with the corresponding atomic forces ($\mathbf{F}_{1}$, $\mathbf{F}_{2}$) are evaluated at each image. The dimer is rotated so as to minimize the total dimer energy, $E$, which is the sum of the energy of the two images, $E=E_{1}+E_{2}$, while keeping the separation between the images fixed. This aligns the orientation of the dimer, $\hat{\mathbf{N}}$, with the minimum mode. An iterative process is used to converge to the direction of the minimum mode to a given tolerance in the force [40]. During the rotation phase, the atomic forces are projected to make them perpendicular to the current dimer orientation, $\mathbf{F}^{\perp}$. Each rotational step involves finding a plane defined by the rotational force and the previous rotation plane, and the rotational force is then zeroed within this plane. Here, a conjugate gradient (CG) method is used where at each step a linear combination of the force vector and the previous search direction is calculated. The coefficient of the previous direction, or the CG parameter, is also a sensitive indicator of convergence. The Polak-Ribiere method [44]:
@@
Once the dimer has been oriented sufficiently close to the minimum mode, a translational step moves the midpoint of the dimer forward. If at least one of the eigenvalues of the Hessian is negative, a modified force, $\mathbf{F}^{\dagger}$, is used, where the component of the force along the dimer direction is inverted: $\mathbf{F}_{R}-2(\mathbf{F}_{R}\cdot\hat{\mathbf{N}})\hat{\mathbf{N}}$. This inversion ensures that the dimer moves uphill on the energy surface along the minimum mode and downhill along all other directions. The dimer is iteratively rotated and translated in this way until the atomic forces become zero to a given tolerance, indicating that a saddle point has been located. The limited memory Broyden-Fletcher-Goldfarb-Shanno algorithm (L-BFGS) on the midpoint of the dimer at each translation step, as it has previously been found to be efficient [29, 47]. The maximum displacement in each iteration is chosen here to be relatively small, 0.05 {Å}, in order to reduce the probability of the dimer search escaping from the saddle point basin of attraction the starting configuration belongs to.
@@ 2.2 Gaussian Process Regression
A Gaussian Process Regression (GPR) is a non-parametric method for learning from data. It can be understood as a generalization of multivariate Gaussian distributions to function spaces, effectively providing a distribution over possible functions that fit the observed data. Here, we use a GPR to learn a relationship between Cartesian coordinates of the atoms and their corresponding energy and atomic forces. The GPR model infers this relationship by finding similarities, determined by means of a covariance function (often referred to as the kernel). The expressivity of these models stems from the fact that they may be understood as the infinite limit of a neural network [38] and that any neural network of a finite width is essentially a GP [50].
@@
where the summations are over all atoms in the system, $(\mathbf{x},\mathbf{x}^{\prime})$ are the input vectors of atomic coordinates, and $r_{i,j}(\mathbf{x})=\sqrt{\sum_{d=1}^{3}(x_{i,d}-x_{j,d})^{2}}$ is the Euclidean distance between atoms $i$ and $j$ within configuration $\mathbf{x}$. $\phi(i,j)$ is a function specifying the type of an atom pair and $l_{\phi(i,j)}$ is the corresponding length scale parameter controlling how quickly the covariance changes as a function of the inverse distance between atoms. The parameter $\sigma_{c}^{2}$ is a constant variance term while $\sigma_{m}^{2}$ is a variance scaling factor for the distance-dependent part of the kernel. This form of the kernel is chosen because the inverse distance term, $1/r_{ij}$, penalizes configurations where atoms are close to each other. This takes into account that atoms cannot occupy the same point in space, effectively modeling the strong repulsive forces at short distances.
@@
The GPR-dimer calculation starts out with up to 6 HF calculations as the dimer is rotated to identify the minimum mode direction, but the usual dimer rotation convergence criteria also apply, so for several systems there are fewer HF calculations in this initial phase. These HF calculations constitute the data set for the initial training of the GPR.
The fitting process involves optimization of the kernel hyperparameters. The computations involve an inversion of a matrix of dimensionality $O(M^{3}N^{3})$ and requires memory on the order of $O(M^{2}N^{2})$, where $M$ is the number of electronic structure calculations that are being used as the training set (i.e. “observations”) and $N$ is the number of atomic coordinates, three times the number of atoms. The hyperparameters are optimized using the scaled conjugate gradient method [36]. A Cholesky decomposition is used to speed up the prediction of multiple data points [45].
@@ 2.3 Computational specifications
The systems analyzed consist of 500 initial configurations of small organic molecules with between 7 and 25 atoms. The electronic structure calculations are carried out at the HF level with a 3-21G basis set using the NWChem [1] software. The spin unrestricted formalism is used for doublets and spin restricted closed shell formalism is used for the singlet configurations. The self-consistent field (SCF) threshold is set at 10-8 Hartree, and the criterion for convergence to a saddle point is that the norm of the atomic force vector drops below 0.01 eV/Å. The electronic structure calculations carried out here are at the same level and use the same software as those of [23].
For the GPR-dimer calculations, which are carried out using EON [6], the geometry verification pre-processing of NWChem is turned off. For both EON and Sella calculations, the ASE [34] file I/O interface to NWChem is used, which means no symmetrization or centering takes place within NWChem, unlike the default settings for NWChem. As mentioned above, the conjugate gradient method [24] is used for the dimer rotations and the limited memory Broyden-Fletcher-Goldfarb-Shanno algorithm [39, 29] is used for the translation. Since the validity of the surrogate energy surfaces is limited to the region where data is available, early stopping criteria are used to prevent the calculation from venturing too far outside the region [31, 30].
@@
The calculations are carried out on the IRHPC elja computer cluster with snakemake as the workflow engine [35]. The cluster nodes have Intel Xeon Platinum 8358 (128 @ 2.600 GHz) CPUs which run Rocky Linux release 8.10. The Hartree Fock calculations are run with NWChem built with spack, and each run requests 16 cores exclusively for NWChem with 3 GB of memory per node. To reduce I/O overheads, the scratch directory was used. dvc was used to vendor intermediate data, and the final csv data required for the plots along with the orgmode literate programming document [12] is part of the associated monorepo.